# New models of nonsingular black hole dark matter from limiting curvature

### v1

@@ V Conclusions
The present work establishes, for some choices of the black hole masses and the regulator function $f(\hat{\ell})$, that the nonsingular black hole metrics we discuss are among the set of possible candidates that might describe primordial black hole dark matter. The present analysis could be generalized by considering other forms for the regulator function $f(\hat{\ell})$, other mass ranges in which the black hole mass would be expected to change appreciably through Hawking radiation over the lifetime of the universe, and other assumptions about the initial black hole mass spectrum. It would also be of interest to see whether a plausible effective description of quantum gravity can be found that leads to metrics of the form we discuss as exact solutions and that might also suggest a preferred form for the regulator function. These topics will be considered in future work.
@@ Appendix A Greybody factors
[…]

### v2

@@ V Conclusions
The present work establishes, for some choices of the black hole masses and the regulator function f(^)f($\hat{\ell}$), that the nonsingular black hole metrics we discuss are among the set of possible candidates that might describe primordial black hole dark matter. The present analysis could be generalized by considering other forms for the regulator function f(^)f($\hat{\ell}$), other mass ranges in which the black hole mass would be expected to change appreciably through Hawking radiation over the lifetime of the universe, and other assumptions about the initial black hole mass spectrum. It would also be of interest to see whether a plausible effective description of quantum gravity can be found that leads to metrics of the form we discuss as exact solutions and that might also suggest a preferred form for the regulator function. These topics will be considered in future work.
@@ Appendix A Greybody factors
$\tau=\left\{\begin{array}[]{ll}\displaystyle 1-3\frac{\hat{\ell}^{2}f(\hat{\ell})}{\hat{r}_{H}^{2}}&\hskip 20.00003pt\mbox{de\,Sitter core,}\\[10.0pt] \displaystyle 1-\frac{\hat{\ell}^{2/3}f(\hat{\ell})}{\hat{r}_{H}}&\hskip 20.00003pt\mbox{Minkowski core.}\end{array}\right$. (44)